Question 1: A $1,010 face value bond is selling in the market place for $934. It matures in 3 years. If keep to maturity, what is the bond's yield to maturity?

Question 2: A banker must earn at least a 2.7% return after expected inflation on short term loans. The inflation rate for the past 6 months has averaged 3.5%. The expected inflation rate for the next twelve months is 7.3%. Nominal interest rates for short term loans were 7.1% last month. What is the minimum nominal interest rate that he should charge for a one year loan?

Question 3: Fred bought a $1000 face value bond issued by Zest Corporation for $1200. The bond matures in 2020 and pays him an annual interest payment of $70. What is Fred's current yield with this bond?

Question 4: A year ago, you purchased a $1000 face value bond for $1,006. A year later you sold the bond for $989 after receiving a coupon payment of $49. What was your rate of capital gain?

Question 5: You need to determine the market value of a $5,000 face value bond maturing in 5 years. The market yield for this type of bond is 4.7%. What is its market value?
Answer #1

(Question 1)

If yield-to-maturity (YTM) is R%, then

Current price x (1 + R)Years to maturity = Face value

$934 x (1 + R)3 = $1,010

(1 + R)3 = $1,010 / $934

(1 + R)3 = 1.0814

Taking cube root,

1 + R = 1.0264

R = 0.0264

R = 2.64%

NOTE: As per Answering Policy, 1st question is answered.